geerky42 (geerky42):

Challenge: Given that \(\Large a^x + a^y = a^z\), where \(\Large a \neq 0\) Find z in term of x and y.

OpenStudy (anonymous):

Good question. Maybe these guys can help you? @UnkleRhaukus @.Sam. @AravindG @ganeshie8 @RadEn @goformit100 @terenzreignz @Preetha @inkyvoyd @nubeer @genius12

OpenStudy (aravindg):

take log on both sides

sam (.sam.):

Does this count? \[\frac{\ln(a^x+a^y)}{lna}=z\]

OpenStudy (aravindg):

:D ^^

OpenStudy (anonymous):

I hope he don't mind a.

terenzreignz (terenzreignz):

^I wanted to do that too, but it seemed too good to be true :(

OpenStudy (unklerhaukus):

x=y=z=2

OpenStudy (aravindg):

\[2a^2=a^2 ??\]

OpenStudy (aravindg):

1=2 ?!

OpenStudy (anonymous):

look like someone broke math.

OpenStudy (unklerhaukus):

does a=a?

OpenStudy (anonymous):

why shouldnt it?

OpenStudy (unklerhaukus):

i would like to withdraw my attempt

OpenStudy (aravindg):

np :)

OpenStudy (unklerhaukus):

how about a=2, x=1, y=1, z=2 z=x+y

OpenStudy (aravindg):

Why are you going for specific cases here @UnkleRhaukus ?

OpenStudy (aravindg):

"Find z in term of x and y". How could constants come as the answer?

OpenStudy (anonymous):

Fermat's Last thm

sam (.sam.):

=.=

OpenStudy (anonymous):

and Beal's conjecture (unproved thm)

OpenStudy (perl):

whats Beal's conjectur

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Beal's_conjecture

OpenStudy (anonymous):

So it's like Fermat's Last Theorem, except that all bases are same instead of exponents, like other way around, right?

OpenStudy (anonymous):

yes. So, can anybody invite Andrew Weyl to Open Study to answer this?

OpenStudy (anonymous):

umm who's Andrew Weyl?

OpenStudy (unklerhaukus):

http://en.wikipedia.org/wiki/Andrew_Wiles

OpenStudy (anonymous):

Oh... haha